OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1+4+7+...+(3n-2)=n(3n-1/2. Answer for Sk: (3k-2)=k(3k-1)/2 Answer for Sk+1: (3(k+1)-2)=(k+1)(3(k+1)-1/2 simplified further: (3k+1)=3k^2+5k+2/2 Please let me know if this is correct. Thank you.

OpenStudy (anonymous):

Sn is actually: 1+4+7+...+(3n-2)=n(3n-1)/2

OpenStudy (anonymous):

\[1+4+7+...+(3n-2)=\frac{n(3n-1)}{2}\] correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Hi Satellite. Thanks for helping again

OpenStudy (anonymous):

I'm getting the hang of the statements with the help. Thanks again!

OpenStudy (anonymous):

i didn't really do anything, just rewrote the statement so i could read it do you know how to replace \(n\) by \(k+1\) ?

OpenStudy (anonymous):

Yes. Thanks. (3(k+1)-2)=(k+1)(3(k+1)-1)/2 simplified further: (3k+1)=3k^2+5k+2/2